OpenStudy (anonymous):

Find an equation, in slope-intercept form, that passes through the point (-4,3) with slope -3. Question 2 options: y= -3x+9 y= -3x-9 y= 3x-9 y = 3x + 9

OpenStudy (anonymous):

y int + -9

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

y-3=-3(x-(-4)) y = -3x-12+3 y =-3x - 9

OpenStudy (anonymous):

tk